# Unleashing the Power of Meta-tuning for Few-shot Generalization Through Sparse Interpolated Experts

### v1

@@ Abstract
Conventional wisdom suggests parameter-efficient fine-tuning of foundation models as the state-of-the-art method for transfer learning in vision, replacing the rich literature of alternatives such as meta-learning. In trying to harness the best of both worlds, meta-tuning introduces a subsequent optimization stage of foundation models but has so far only shown limited success and crucially tends to underperform on out-of-domain (OOD) tasks. In this paper, we introduce Sparse MetA-Tuning (SMAT), a method inspired by sparse mixture-of-experts approaches and trained to isolate subsets of pre-trained parameters automatically for meta-tuning on each task. SMAT successfully overcomes OOD sensitivity and delivers on the promise of enhancing the transfer abilities of vision foundation models beyond parameter-efficient finetuning. We establish new state-of-the-art results on a challenging combination of Meta-Dataset augmented with additional OOD tasks in both zero-shot and gradient-based adaptation settings. In addition, we provide a thorough analysis of the superiority of learned over hand-designed sparsity patterns for sparse expert methods and the pivotal importance of the sparsity level in balancing between in-domain and out-of-domain generalization. Our code is publicly available.
@@ 1 Introduction
Figure 1: Average testing accuracy during meta-training for meta-tuning methods. SMAT yields better in-domain (ID) and out-of-domain (OOD) results and shows an attractive learning speedup.
@@
Despite initial progress made, the exploration of meta-tuning in vision still remains notably limited to date. Hu et al. propose a pre-training $\rightarrow$ meta-training $\rightarrow$ fine-tuning pipeline, dubbed PMF, for enhancing the few-shot learning performance of the resulting model relative to the default pre-training $\rightarrow$ fine-tuning approach. With the design principle of simplicity in mind, PMF meta-trains all parameters in a vision transformer using Prototypical Networks (Snell et al., 2017) starting from a pre-trained initialization, yielding the state-of-the-art performance on popular meta-learning benchmarks such as Meta-Dataset (Triantafillou et al., 2020). Despite such promising reported results, we find that this intuitive approach tends to underperform on downstream few-shot tasks, particularly when testing for out-of-domain (OOD) tasks (i.e., tasks dissimilar to the ones presented during the meta-training stage).
We hypothesize that this low generalization performance on OOD tasks stems from two major factors. (i) The strong emphasis on learning from small amounts of data using a limited number of optimization steps in meta-learning can lead to algorithms that are “greedy” wrt. to the distribution of tasks presented, sacrificing more generalizable features for performance on the distribution at hand. This leads to a risk of meta-overfitting, a phenomenon previously observed (e.g. Zintgraf et al., 2019; Yao et al., 2021; Chen et al., 2022). (ii) When meta-tuning tasks are diverse, the default setting of updating all parameters suffers from task interference, making optimization unstable and thereby reducing generalization performance.
@@
2 Preliminary: Meta-tuning
Meta-tuning aims to improve the few-show learning performance of a pre-trained model on downstream testing task episodes - usually by directly optimizing the few-shot learning objective over a collection of labeled training task episodes  (Hu et al., 2022; Min et al., 2021).
Given a powerful pre-trained backbone $f_{\bm{\theta}^{\mathtt{pre}}}$, we assume the availability of a task distribution ${\mathcal{P}}_{ID}({\mathcal{T}})$ from which we may sample fully labeled training task episodes ${\mathcal{T}}_{i}\sim{\mathcal{P}}_{ID}({\mathcal{T}})$. Note that we explicitly denote this as in-distribution ($ID$). In particular, in the supervised setting, each training task consists of a fully labeled support ${\mathcal{T}}_{i}^{s}\!:=\!\{\mathbf{x}_{i_{j}}^{s},\mathbf{y}_{i,j}^{s}\}_{j}^{N_{i}^{s}}$ and query ${\mathcal{T}}_{i}^{s}\!:=\!\{\mathbf{x}_{i_{j}}^{q},\mathbf{y}_{i,j}^{q}\}_{j}^{N_{i}^{q}}$ set of $N_{i}^{s}$, $N_{i}^{q}$ input-label pairs, respectively. Meta-tuning is then realized through the typical episodic-learning setting familiar from meta-learning, i.e. the direct optimization of $\bm{\theta}$ on the few-shot learning objective, i.e., $\operatorname*{arg\,min}_{\Theta}\mathbb{E}_{{\mathcal{P}}_{ID}}[{\mathcal{L}}_{i}(f_{\Theta}(\mathbf{X}^{q}_{i},{\mathcal{T}}^{s}_{i}),\mathbf{Y}^{q}_{i})]$ over task episodes sampled from ${\mathcal{P}}_{ID}({\mathcal{T}})$. At test time, we expect to encounter both ID testing tasks i.e., $\tilde{{\mathcal{T}}}:=\{\tilde{{\mathcal{T}}}^{s}_{i},\tilde{\mathbf{X}}_{i}^{q}\}\sim{\mathcal{P}}_{ID}$, and unseen OOD testing tasks which come from a different task distribution ${\mathcal{P}}_{OOD}\neq{\mathcal{P}}_{ID}$. Our goal is therefore to achieve the highest generalization performance on the query sets of both ID and OOD downstream testing tasks.
Figure 2: Overview of the proposed Sparse MetA-Tuning approach, showing the meta-training and inference for one particular task ${\mathcal{T}}_{i}$.
3 SMAT: Sparse MetA-Tuning
[…]
As discussed in Section 1, naively sharing and updating all pre-trained parameters across all tasks in meta-tuning leads to task interference in optimization. To address this issue, we instead hypothesize that the solution for each task (ID or OOD) comprises a task-specific mixture of a common pool of knowledge or skills covering a broad range of tasks. The knowledge pool is represented by distinct sets of model parameters (i.e., experts), which can be combined cooperatively as a complementary addition to the pre-trained model to promote systematic generalization. Formally, we assume that each task-specific model $\bm{\theta}_{i}$ is derived from aggregating the experts via a task-specific weighted sum in the parameter space:
where ${\bm{\theta}}_{m}^{\delta}$ represents the $m$-th expert in the pool ${\mathcal{M}}$, and $\bm{\alpha}_{i}:=[\alpha_{i,1},\alpha_{i,2},...,\alpha_{i,|{\mathcal{M}}|}]$ are the weights. This way of merging experts in Eqn (1) is appealing due to its high expressiveness, as supported by recent findings revealing that merging multiple sets of specialized parameters through simple arithmetic can lead to a better overall multi-task solution (Ilharco et al., 2023; Matena & Raffel, 2021). Moreover, the complementary addition of experts to the pre-trained backbone effectively decouples the two for optimization - meta-training the experts while keeping $\bm{\theta}^{\mathtt{pre}}$ frozen provides sufficient model capacity while preserving all pre-trained knowledge.
That said, meta-tuning, in our case, is seen as a procedure for discovering a generalizable selection rule for assigning appropriate experts to tasks.
Sparsification of experts.    At the core of our method lies the question on where and how to find a set of experts in the model during meta-tuning. Previous studies on partitioned meta-learning, as well as recent works on Mixture-of-Experts, propose a seemingly reasonable solution: Experts are incorporated into a heuristically (and a-priori) chosen subset of modules in the model. For example, the batch-norm (BN) layers in ResNets (Triantafillou et al., 2021b) or the MLP layers in Vision Transformers (ViT) (Puigcerver et al., 2023). Although these approaches based on fixed partitioning are more memory-efficient and generally result in improved performance compared to their non-expert counterparts, they suffer from a significant bias due to their manually crafted selection (i.e., the concrete choice of $\bm{\theta}_{m}$). This bias may be suboptimal when it conflicts with the intricate dynamics of meta-learning and, in addition, is not always model-agnostic, e.g., since there are no BN layers in ViT.
Thus, instead of specifying experts as prior knowledge, we propose to automatically identify the sparsity patterns in experts through meta-tuning using a maximum likelihood objective while considering sparsity constraints on the experts’ capacity, thus encouraging specialization. The meta-learned sparsity patterns not only exhibit minimal bias but also allow differentiation among experts’ architecture - which is essential for addressing all various kinds of distribution shifts (Lee et al., 2022).
Specifically, we induce a learnable sparsity pattern in the $m$-th expert in the form of a sparse reparametermization $\bm{\theta}_{\delta}\odot\bm{z}_{m}$ of the dense expert, where $\bm{z}_{m}$ denotes a learnable binary mask on which we impose the sparsity constraint. To learn discrete binary masks through gradient optimization, we take inspiration from (Schwarz & Teh, 2022), and reparameterize the $m$-th binary mask using an underlying continuous distribution $q_{\bm{\phi}_{m}}$ governed by parameters $\bm{\phi}_{m}$. The reparameterization samples, i.e., $\mathbf{s}_{m}\!\sim\!q_{\bm{\phi}_{m}}(\mathbf{s})$, can be transformed to have values exactly 00 and $1$ through a hard rectification $\mathbf{z}_{m}\!=\!g(\mathbf{s}_{m})\!:=\!\min(1,\max(0,\mathbf{s}_{m}))$. As a result, sparsity in masks can be enforced by limiting the probability of $s_{m}$ being non-zero, which can be straight-forwardly expressed through the CDF $Q$ of $q_{\bm{\phi}_{m}}$, i.e. $1-Q_{\bm{\phi}_{m}}(\mathbf{s}\leq 0)$. Choosing $q_{\bm{\phi}_{m}}$ as the stretched hard concrete distribution (Louizos et al., 2018) enables both effective gradient-based learning through reparameterization as well as analytical evaluation of the CDF.
From sparse experts to sparse interpolated experts.    Though the aforementioned sparse reparameterization eventually reduces the final parameter counts by removing zeros in expert parameters, it also leads to a strong increase in the number of parameters compared to its non-sparse counterpart at the beginning of meta-tuning – due to learning $\bm{z}_{m}$ and $\bm{\theta}_{m}^{\delta}$ concurrently. Moreover, assigning each expert its own dense underlying parameters $\bm{\theta}_{m}^{\delta}$ greatly hinders knowledge transfer among experts, which contradicts the principles of partitioned meta-learning emphasizing collaboration of both task-agnostic and task-specific components.
To tackle these issues, we propose sharing dense parameters across sparse reparameterization of different experts, i.e., $\bm{\theta}_{m}^{\delta}=\bm{\theta}^{\delta},\forall m$. By rearranging Eq. 1 slightly, the experts now essentially become different sparse interpolations between the same pre-trained and meta-tuned checkpoints. Although this formulation of MoE may initially appear bold, it is well-supported by recent works (Panigrahi et al., 2023b) which suggest that multiple task-specific optimal points can coexist between the same set of pre-trained and meta-tuned models, offering favorable trade-offs for both in-distribution (ID) and some out-of-distribution (OOD) performance. Therefore, our approach of learning sparse interpolated experts can be seen as an inductive bias that promotes the recovery of these optimal interpolation points through meta-tuning. Favorably, this partitioning allows for knowledge transfer among experts through the overlapping regions in their masks.
Meta-learning expert selection through a hypernetwork.     In theory, there are two possible approaches to achieve task-specific inference for the expert merging scores $\bm{\alpha}_{i}$. The first one is a meta-learned merging score initialization $\bm{\alpha}_{0}$ combined with inner-loop gradient-descent on the task support set; the second involves a meta-learned global hypernetwork that directly outputs $\bm{\alpha}_{i}$ conditioned on the task support set ${\mathcal{T}}_{i}^{s}$. We opt for the latter approach due to its better scalability with large-scale models. For implementation, we use the pre-trained model $f_{\bm{\theta}^{\mathtt{pre}}}$ to encode each support set image $\mathbf{x}_{i,j}^{s}\in\mathcal{T}_{i}^{s}$ into a vector embedding $f_{\bm{\theta}^{\mathtt{pre}}}(\mathbf{x}_{i,j}^{s})$ at the output of its last layer. These embeddings are then aggregated into class prototypes. We concatenate the prototypes into a sequence and feed it into a single trainable transformer block to obtain $\bm{\alpha}_{i}^{\prime}\in\mathbb{R}^{{\mathcal{M}}}$ as the output. We treat $\bm{\alpha}_{i}^{\prime}$ as the logits for the activation of each expert. Instead of choosing the top-k activated ones, which can cause training instability issues and restrict the number of experts per task, we employ the Gumbel-Sigmoid trick (Jang et al., 2017) to sample a soft activation value $\in(0,1)$ for each mask, followed by normalization to obtain $\bm{\alpha}_{i}\in(0,1)^{|{\mathcal{M}}|}$.
Enhancing expert specialization through task-specific dense teachers.    Specialization and cooperation among experts play a crucial role in a MoE model. One way to promote specialization is by penalizing the similarity between experts. For example, an orthogonal penalty can be applied to pairs of experts. However, incorporating such explicit penalties makes the optimization problem in the meta-objective more challenging, as it introduces a trade-off with respect to the few-shot prediction performance. To circumvent this trade-off, we propose an alternative approach that utilizes a knowledge distillation loss $\mathcal{L}^{\mathtt{kd}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i},\bm{\theta}_{i}^{\mathtt{tr}})$ (Hinton et al., 2015) between the merged MoE model $\bm{\theta}_{i}$ and a teacher network $\bm{\theta}_{i}^{\mathtt{tr}}$. By using a highly task-specific teacher, the distillation loss imposes the weighted-sum of experts that constitute each $\bm{\theta}_{i}$ to mimic the behaviour of the teacher, thus enhancing both specialization and cooperation implicitly through knowledge transfer. Furthermore, the predictive performance of $\bm{\theta}_{i}$ is not impeded because the loss consistently encourages the current student to achieve a better task-specific generalization on the query set, similar to that of the teacher.
We generate the teacher $\bm{\theta}_{i}^{\mathtt{tr}}$ dynamically in each meta-training episode by performing $K$-step gradient descent starting from $\bm{\theta}_{i}$ on the query loss $\mathcal{L}^{\mathtt{ce}}_{\mathcal{T}_{i}^{q}}$ (we find $K=1$ is sufficient in practice). It is worth noting that unlike $\bm{\theta}_{i}$, we do not limit the capacity of $\bm{\theta}_{i}^{\mathtt{tr}}$ through sparse regularization. This results in a teacher model that is both dense in modulation (i.e., $\bm{\theta}_{i}^{\mathtt{tr}}-\bm{\theta}^{\mathtt{pre}}$), making it expressive and highly task-specific. Importantly, we do not propagate loss gradients through $\bm{\theta}_{i}^{\mathtt{tr}}$. This approach is thus similar to the use of bootstrapping in optimization-based meta-learning (Flennerhag et al., 2022; Tack et al., 2024).
$\min\limits_{\bm{\theta}^{\delta},\bm{\zeta},\bm{\Phi}}\max\limits_{\bm{\lambda}}$ $\mathbb{E}_{{\mathcal{T}}_{i}\sim p^{\text{train}}({\mathcal{T}})}\big{[}{\mathcal{L}}^{\mathtt{ce}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i})+\mathcal{L}^{\mathtt{kd}}_{\mathcal{T}_{i}^{q}}(\bm{\theta}_{i},\bm{\theta}_{i}^{\mathtt{tr}})\big{]}$
$+\sum_{m=1}^{|{\mathcal{M}}|}\lambda_{m}(\frac{1}{|\bm{\phi}_{m}|}\sum_{k=1}^{|\bm{\phi}_{m}|}\tau-(1-Q_{\bm{\phi}_{m}}(s_{k}\leq 0))),$
where $\bm{\theta}_{i}=\bm{\theta}^{\mathtt{pre}}+\bm{\theta}^{\delta}\odot\sum_{m=1}^{|{\mathcal{M}}|}\alpha_{i,m}\bm{z}_{m},$
$\bm{z}_{m}\sim q_{\bm{\phi}_{m}};\bm{\alpha}_{i}\sim h_{\zeta},$ (2)
where the objective is the minimization problem in the first line while the aforementioned sparsity constraints translate to the maximization problem in the second. We put the sparsity constraints on individual masks, hence experts, as opposed to the overall $\bm{\theta}_{i}$ to allow task-dependent model capacity in $\bm{\theta}_{i}$ through selective merging. For simplicity, we set a common constraint for all masks i.e., $\tau_{m}=\tau,\forall m$, which is treated as a hyperparameter. We use simultaneous gradient descent and projected gradient ascent for optimizing Eqn. (3.1). In order to avoid penalizing the model’s capacity excessively due to exceeding the sparsity constraints, we reset $\lambda_{m}$ to zero once its corresponding sparsity constraint has been satisfied (Gallego-Posada et al., 2022). This results in final sparsity close to the target $\tau$, enabling precise control of the expert sparsity levels. The pseudocodes for meta-training can be found in Section A.1.
3.2 Meta-testing time adaptation
Gradient-free optimization for expert selection.    Prior work has discovered that further task-specific adaptation of a meta-trained model is essential for improving its performance on OOD tasks during meta-testing (Hu et al., 2022; Chen et al., 2023; Li et al., 2022). Although effective, one major limitation of these adaptation methods is the reliance on back-propagation of the gradients, which can be expensive, making these methods inefficient and potentially impractical due to poor scaling with model size. To this end, we propose an adaptation strategy, specifically designed for SMAT, that bypasses the gradient computation. At the core of our method lies the hypothesis that each expert selection score can be descreatzed i.e, either 0 or 1, which aligns with the intuition that each expert knowledge is either required or not for solving each tasks. We can thus optimize the expert selection score in a binary hypothesize space i.e, $\bm{\alpha}_{i}\in\{0,1\}^{{{\mathcal{M}}}}$. A more detailed description of the algorithm can be found in Section A.2.
Gradient-based fine-tuning with SMAT.    Following (Hu et al., 2022), we also consider gradient-based fine-tuning of our meta-tuned model at meta-testing time. With only few changes, our method is fully compatible with existing full fine-tuning (i.e., fine-tuning the entire model), and parameter-efficient fine-tuning (PEFT) techniques. Specifically, we use $\bm{\theta}_{i}$ in Eqn. (3.1) as the task-specific model initialization, before applying any off-the-self fine-tuning technique for further optimizing $\bm{\theta}_{i}$ on the support set of each task. The pseudocode, as well as ablation results on fine-tuning SMAT, at meta-testing time, using our proposed methods are shown in the Appendix.
@@ 4 Related Work
Few-shot learning and meta-tuning.    Much of few-shot learning (FSL) relies on meta-learning across so-called N-way-K-shot meta-training episodes, which can then be utilized for data-efficient learning during (Hospedales et al., 2020). Meta-learned inductive biases may take the form of a model initialization (Finn et al., 2017), a learned metric (Snell et al., 2017), a Bayesian prior (Grant et al., 2018) or an optimization strategy (Li et al., 2017). A particular subdomain of FSL, namely, cross-domain FSL algorithms (Li et al., 2022; Triantafillou et al., 2020; Liu et al., 2021; Bateni et al., 2019), explicitly deals with task-distribution shift between meta-training and -testing. Nevertheless, most architectures used in existing work are limited in scale and without large-scale pre-training. Transitioning into the LLM era,  (Min et al., 2021) first study meta-training a pre-trained LLM on a large collection of few-shot in-context learning tasks. Their results highlight the effectiveness of meta-training on improving few-shot in-context learning generalization of powerful pre-trained transformers; motivating several follow-up works in the field of meta-tuning in NLP (Gao et al., 2021; Min et al., 2021; Chen et al., 2021). In computer vision,  Hu et al. propose the simple pre-training, meta-training then fine-tuning (PMF) pipeline and achieve SOTA performance (2022).
@@
Since then, an increasing amount of attention has been placed on adapters, i.e. compact, parameter-efficient modules which have been shown to be particularly impactful when fine-tuning foundation models, thus being particularly suitable for Meta-Tuning. Popular adapter types such as FiLM (Perez et al., 2018), LoRA (Hu et al., 2021) as well as various alternatives feature in various meta-learning and episodic-training methods (e.g. Requeima et al., 2019; Triantafillou et al., 2021a; Shysheya et al., 2022; Schwarz et al., 2023).

### v3

@@ Abstract
Recent successes suggest that parameter-efficient fine-tuning of foundation models is becoming the state-of-the-art method for transfer learning in vision, gradually replacing the rich literature of alternatives such as meta-learning. In trying to harness the best of both worlds, meta-tuning introduces a subsequent optimization stage of foundation models but has so far only shown limited success and crucially tends to underperform on out-of-distribution (OOD) tasks. In this paper, we introduce Sparse MetA-Tuning (SMAT), a method inspired by sparse mixture-of-experts approaches and trained to isolate subsets of pre-trained parameters automatically for meta-tuning on each task. SMAT successfully overcomes OOD sensitivity and delivers on the promise of enhancing the transfer abilities of vision foundation models beyond parameter-efficient fine-tuning. We establish new state-of-the-art results on a challenging combination of Meta-Dataset augmented with additional OOD tasks in both zero-shot and gradient-based adaptation settings. In addition, we provide a thorough analysis of the superiority of learned over hand-designed sparsity patterns for sparse expert methods and the pivotal importance of the sparsity level in balancing between in-distribution and out-of-distribution generalization. Our code and models are publicly available.
@@ 1 Introduction
Figure 1: Average testing accuracy during meta-training for meta-tuning methods. SMAT yields better ID and OOD results and shows an attractive learning speedup.
@@
Despite initial progress made, the exploration of meta-tuning in vision still remains notably limited to date. Hu et al. propose a pre-training $\rightarrow$ meta-training $\rightarrow$ fine-tuning pipeline, dubbed PMF (2022), for enhancing the few-shot learning performance of the resulting model relative to the default pre-training $\rightarrow$ fine-tuning approach. With the design principle of simplicity in mind, PMF meta-trains all parameters in a vision transformer using Prototypical Networks (Snell et al., 2017) starting from a pre-trained initialization, yielding the state-of-the-art performance on popular meta-learning benchmarks such as Meta-Dataset (Triantafillou et al., 2020). Despite such promising reported results, we find that this intuitive approach tends to underperform on downstream few-shot tasks, particularly when testing for out-of-distribution (OOD) tasks (i.e., tasks dissimilar to the ones presented during the meta-training stage).
We hypothesize that this low generalization performance on OOD tasks stems from two major factors. (i) The strong emphasis on learning from small amounts of data using a limited number of optimization steps in meta-learning can lead to algorithms that are “greedy” w.r.t. the distribution of tasks presented, sacrificing more generalizable features for performance on the distribution at hand. This leads to a risk of meta-overfitting, a phenomenon previously observed (e.g. Zintgraf et al., 2019; Yao et al., 2021; Chen et al., 2022). (ii) When meta-tuning tasks are diverse, the default setting of updating all parameters suffers from task interference, making optimization unstable and thereby reducing generalization performance.
@@ 2 Related Work
Few-shot learning and meta-tuning.    Much of few-shot learning (FSL) relies on extracting transferable prior knowledge from a collection of few-shot training task episodes through meta-learning (Hospedales et al., 2020), which can then be utilized for data-efficient learning on unseen but related downstream FSL tasks at test time. Meta-learned inductive biases may take the form of a model initialization (Finn et al., 2017), a learned metric (Snell et al., 2017), a Bayesian prior (Grant et al., 2018) or an optimization strategy (Li et al., 2017). A particular subdomain of FSL, namely, cross-domain FSL algorithms (Li et al., 2022; Triantafillou et al., 2020; Liu et al., 2021; Bateni et al., 2019), explicitly deals with task-distribution shift between meta-training and -testing. Nevertheless, most architectures used in existing work are limited in scale and without large-scale pre-training. Transitioning into the LLM era,  (Min et al., 2021) first study meta-training a pre-trained LLM on a large collection of few-shot in-context learning tasks. Their results highlight the effectiveness of meta-training on improving few-shot in-context learning generalization of powerful pre-trained transformers; motivating several follow-up works in the field of meta-tuning in NLP (Gao et al., 2021; Min et al., 2021; Chen et al., 2021). In computer vision,  Hu et al. propose the simple pre-training, meta-training then fine-tuning (PMF) pipeline (2022) and achieve SOTA performance. Concurrent to our work, Eustratiadis et al. explore meta-tuning from an orthogonal direction to ours by proposing a neural architecture search algorithm (2024) designed to find the optimal model configuration (e.g., optimal arrangement of adapters) for fine-tuning a meta-tuned model for downstream adaptation.
@@
Since then, an increasing amount of attention has been placed on adapters, i.e. compact, parameter-efficient modules which have been shown to be particularly impactful when fine-tuning foundation models, thus being particularly suitable for Meta-Tuning. Popular adapter types such as FiLM (Perez et al., 2018), LoRA (Hu et al., 2021) as well as various alternatives feature in various episodic-training methods (e.g. Requeima et al., 2019; Triantafillou et al., 2021a; Shysheya et al., 2022; Schwarz et al., 2023).